OpenStudy (anonymous):

The length of the hypotenuse of a right triangle is 13 m. The length of one leg is 5 cm. Find the length of the other leg.

OpenStudy (aripotta):

if you're only 5, why do you need to do this. are you even in school yet?

OpenStudy (anonymous):

hehehe im trying to get smarter.........wats wrong with that?

OpenStudy (aripotta):

..right.

OpenStudy (anonymous):

hehehehe

OpenStudy (anonymous):

right well if you want to be smarter learn how to transpose equations. The equation for any length of triangle is hypotenuse law. A^2+B^2=C^2 for triangle ABC C being the hypotenuse.

OpenStudy (anonymous):

okay soo the answer would beeeeeeeeeeeeeeeeeeeee 224?

OpenStudy (anonymous):

no, look

OpenStudy (anonymous):

A^2 =? B^2= 5 C^2=13 So A^2+5^2=13^2 So A^2=13^2-5^2 Therefore A= sqrt(13^2-5^2)

OpenStudy (anonymous):

oh that makes snce

OpenStudy (anonymous):

Glad to help, so you should get A =12